# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 5.1. Rigidity of the mixed discrete conformal structure $\mathrm{I}$
Since $u_{j},u_{k}\in\mathbb{R}_{<0}$, then $\eta_{jk}>0$ and $u_{j}+u_{k}>-\sqrt{2\eta_{jk}}$. Similarly,
$\zeta_{ij}=\frac{\cosh l_{ij}-\cosh(f_{j}-f_{i})}{e^{f_{i}+f_{j}}}>\frac{1-\frac{1}{2}(e^{f_{j}-f_{i}}+e^{f_{i}-f_{j}})}{e^{f_{i}+f_{j}}}=-\frac{1}{2}(e^{-f_{i}}-e^{-f_{j}})^{2}=-\frac{1}{2}(u_{i}+u_{j})^{2},$
which implies $(u_{i}+u_{j})^{2}>-2\zeta_{ij}$. Moreover, if $\zeta_{ij}>0$, then $u_{i}+u_{j}\in\mathbb{R}$; if $\zeta_{ij}<0$, then $u_{i}+u_{j}>\sqrt{-2\zeta_{ij}}$ or $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}$. To remove the solution $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}$, we add the condition that $\eta_{jk}+\zeta_{ij}\leq 0$. In fact, $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}\leq-\sqrt{2\eta_{jk}}<u_{j}+u_{k}$, which contradicts with the fact $u_{i}>0$ and $u_{k}<0$. Similar arguments are also suitable for $\zeta_{ik}$. Hence, we have the following theorem, which characterizes the admissible space of $u$.
@@
Suppose $(\Sigma,\mathcal{T},\zeta,\eta)$ is a weighted triangulated surface with boundary, and the weights $\zeta\in\mathbb{R}^{E}$ and $\eta\in\mathbb{R}_{>0}^{E}$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (43). If one of the following conditions is satisfied
@@
(44) $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=$ $u(\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik}))$
$=$ $\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},$
$\ \ (u_{i}+u_{j})^{2}>-2\zeta_{ij},(u_{i}+u_{k})^{2}>-2\zeta_{ik}\}$
of $u$ is a convex polytope. As a result, the admissible space
$\mathcal{U}(\zeta,\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$
is also a convex polytope.
@@
According to the arguments above Theorem 5.1, we can describe the admissible space $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ explicitly. For the case $\mathrm{(i)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}}\}.$
For the case $\mathrm{(ii)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{k}>\sqrt{-2\zeta_{ik}}\}.$
For the case $\mathrm{(iii)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\zeta_{ij}}\}.$
For the case $\mathrm{(iv)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|$ $u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\zeta_{ij}},$
$u_{i}+u_{k}>\sqrt{-2\zeta_{ik}}\}.$
Each of these four cases shows that $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ is a convex polytope. Q.E.D.
@@
Under the same assumptions as in Theorem 5.1. The Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite on $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$. As a result, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\zeta,\eta)$.
@@
The formula (30) implies
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{5}.$
Then
where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Moreover, $\det Q_{1}>0$.
Since $d_{ij}<0,d_{ik}<0$ and the others are positive for any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), then
@@
where (1) is used in the second line. Thus
(45) $\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{5}<0,$
which implies the matrix $\Lambda_{ijk}$ is non-degenerate. The symmetry of the matrix $\Lambda_{ijk}$ follows from Lemma 2.22. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}$. Then $a_{iij}<0,\ a_{iik}<0,\ a_{jjk}>0,\ a_{kkj}>0,\ a_{jji}>0,\ a_{kki}>0$. Define
$Q:=Q_{1}Q_{2}Q_{5}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ -a_{iik}&0&a_{kki}\\ -a_{iij}&a_{jji}&0\\ \end{array}\right).$
By similar arguments in the proof of Theorem 3.2, we will show that the determinants of the $1\times 1$ and $2\times 2$ order principal minor of $Q$ are positive. Let $m_{rs}$ be the entry of $Q$ at $r$-th row and $s$-th column. Then
@@
Then the determinants of the $1\times 1$ principal submatrices of $Q$ is $m_{11}>0$ by $a_{iik}<0,\ a_{iij}<0$. And the determinants of the $2\times 2$ principal submatrices of $Q$ is
@@
The following theorem gives the rigidity of the mixed discrete conformal structure $\mathrm{I}$, which generalizes the case $\mathrm{(iii)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it here.
@@
Under the same assumptions as in Theorem 5.1. The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}_{>0}^{E}$. In particular, the map from $f$ to $K$ is a smooth embedding.
5.2. Existence of the mixed discrete conformal structure $\mathrm{I}$
This proof is similar to that of Theorem 3.5 for the discrete conformal structure (A1).
@@ 5.2. Existence of the mixed discrete conformal structure $\mathrm{I}$
Under the same assumptions as in Theorem 5.1. The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
To prove Theorem 5.4, we need the following two lemmas.
@@
Under the same assumptions as in Theorem 5.1. For any constants $c_{1},...,c_{6}\in\mathbb{R}$, if one of the following conditions is satisfied
@@
$\lim f_{i}=-\infty,\ \lim f_{j}=c_{5},\ \lim f_{k}=c_{6}$.
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
@@
By (43), we have
$\cosh l_{ij}$ $=e^{f_{i}+f_{j}}(\zeta_{ij}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{j}}),$
$\cosh l_{ik}$ $=e^{f_{i}+f_{k}}(\zeta_{ik}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{k}}),$
@@
For the case $\mathrm{(i)}$, if $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law,
@@
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs of the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$ are similar, so we omit them here.
For the case $\mathrm{(iv)}$, if $\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (43), we have
$\cosh l_{ij}=e^{f_{j}-f_{i}}(\zeta_{ij}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{j}}+\frac{1}{2}),$
$\cosh l_{ik}=e^{f_{k}-f_{i}}(\zeta_{ik}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{k}}+\frac{1}{2}),$
@@
Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}},\ \lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$ and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants. By the cosine law,
@@
which implies $\theta^{jk}_{i}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(v)}$ and $\mathrm{(vi)}$, and we omit them here. Q.E.D.
@@
Under the same assumptions as in Theorem 5.1. For any constants $c_{1},...,c_{5}\in\mathbb{R}$, if one of the following conditions is satisfied
@@
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{4},\ \lim f_{k}=c_{5}$.
Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
@@
For the case $\mathrm{(i)}$, if $\lim f_{j}=+\infty,\ \lim f_{i}=+\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law,
@@
which implies $\theta^{ik}_{j}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(ii)}$, $\mathrm{(v)}$ and $\mathrm{(vi)}$, and we omit them.
For the case $\mathrm{(iii)}$, if $\lim f_{j}=+\infty,\ \lim f_{i}=-\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$$\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$ and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants. By the cosine law,
@@
which implies $\theta^{ik}_{j}\rightarrow 0$. The same proof is suitable for the case $\mathrm{(iv)}$, so we omit it. Q.E.D.
Proof of Theorem 5.4: Let the image of $K$ is $X$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the injectivity of the map $f$. To prove the closeness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $f^{(m)}$ in $\Omega(\zeta,\eta)=\cap_{\{ijk\}\in F}\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\zeta,\eta)$.
Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\zeta,\eta)$. The boundary of the admissible space $\Omega(\zeta,\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
@@
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$.
@@
There exists $i\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), such that $f_{i}^{(m)}\rightarrow\pm\infty$.

### v3

@@ 6.1. Rigidity of the mixed discrete conformal structure III
Since $u_{j}\in\mathbb{R}_{<0}$ and $u_{k}\in\mathbb{R}_{<0}$, it follows that $\eta_{jk}>0$ and $u_{j}+u_{k}>-\sqrt{2\eta_{jk}}$. Similarly,
$\eta_{ij}=\frac{\cosh l_{ij}-\cosh(f_{j}-f_{i})}{e^{f_{i}+f_{j}}}>\frac{1-\frac{1}{2}(e^{f_{j}-f_{i}}+e^{f_{i}-f_{j}})}{e^{f_{i}+f_{j}}}=-\frac{1}{2}(e^{-f_{i}}-e^{-f_{j}})^{2}=-\frac{1}{2}(u_{i}+u_{j})^{2},$
which implies $(u_{i}+u_{j})^{2}>-2\eta_{ij}$. Moreover, if $\eta_{ij}>0$, then $u_{i}+u_{j}\in\mathbb{R}$; if $\eta_{ij}\leq 0$, then $u_{i}+u_{j}>\sqrt{-2\eta_{ij}}$ or $u_{i}+u_{j}<-\sqrt{-2\eta_{ij}}$. To exclude the solution $u_{i}+u_{j}<-\sqrt{-2\eta_{ij}}$, we add the condition that $\eta_{jk}+\eta_{ij}\leq 0$. In fact, $u_{i}+u_{j}<-\sqrt{-2\eta_{ij}}\leq-\sqrt{2\eta_{jk}}<u_{j}+u_{k}$, which contradicts the fact $u_{i}\in\mathbb{R}_{>0}$ and $u_{k}\in\mathbb{R}_{<0}$. Similar arguments also apply for $\eta_{ik}$. Consequently, we obtain the following theorem, which characterizes the admissible space of $u$.
@@
Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (48). If one of the following conditions is satisfied
@@
(49) $\mathcal{U}_{ijk}(\eta)$ $=u(\Omega_{ijk}(\eta))$
=$=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}},$
$\ \ \ \ \ \ \ \ (u_{i}+u_{j})^{2}>-2\eta_{ij},(u_{i}+u_{k})^{2}>-2\eta_{ik}\}$
of $u$ is a convex polytope.
@@
According to the arguments above Theorem 6.1, we can explicitly characterize the admissible space $\mathcal{U}_{ijk}(\eta)$. For the case (i), the admissible space (49) reduces to
$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}}\}.$
For the case (ii), the admissible space (49) reduces to
$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{k}>\sqrt{-2\eta_{ik}}\}.$
For the case (iii), the admissible space (49) reduces to
$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\eta_{ij}}\}.$
For the case (iv), the admissible space (49) reduces to
$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid$ $\ u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\eta_{ij}},$
$\ u_{i}+u_{k}>\sqrt{-2\eta_{ik}}\}.$
Each of these four cases shows that $\mathcal{U}_{ijk}(\eta)$ is a convex polytope. Q.E.D.
Corollary 6.2.
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary. If one of the following conditions holds:
(1):
All right-angled hyperbolic hexagons are Type-III (i.e., the edge lengths of any right-angled hyperbolic hexagons are given by (48)), and the weight $\eta$ satisfies one of the conditions (i)–(iv) in Theorem 6.1;
(2):
Not all right-angled hyperbolic hexagons are of Type-III (i.e., there exists at least one right-angled hyperbolic hexagon with edge lengths given by (7)). In this case, the weight $\eta$ satisfies one of the conditions (i)–(iv) in Theorem 6.1 for those right-angled hyperbolic hexagons with edge lengths given by (48), and $\eta\in\mathbb{R}_{>0}^{E}$ for those with edge lengths given by (7).
Then the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is a convex polytope on $(\Sigma,\mathcal{T},\eta)$,
Proof.
The case (1) follows from Theorem 6.1. The case (2) follows from (47) and Theorem 6.1. Indeed, by (47), the admissible space $\mathcal{U}_{ijk}(\eta)$ of a right-angled hyperbolic hexagon $\{ijk\}$ with edge lengths given by (7) is a convex polytope. Q.E.D.
@@
Under the same assumptions as those in Theorem 6.1, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite on $\mathcal{U}_{ijk}(\eta)$.
@@
For simplicity, we adopt the notations introduced in Subsection 3.2. Formula (29) implies
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{6}.$
Then $\Lambda_{ijk}$ can be expressed as
where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively. Furthermore, $\det Q_{1}>0$.
For any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), since $d_{ij}<0$, $d_{ik}<0$, $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$, it follows that
@@
where (2) is used in the second line. Consequently,
$\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{6}<0,$
which implies the matrix $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}$. Then $a_{iij}<0$, $a_{iik}<0$, $a_{jjk}>0$, $a_{kkj}>0$, $a_{jji}>0$, and $a_{kki}>0$. Define
$Q:=Q_{1}Q_{2}Q_{6}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ -a_{iik}&0&a_{kki}\\ -a_{iij}&a_{jji}&0\\ \end{array}\right).$
By arguments analogous to those in the proof of Theorem 3.2, we show that the $1\times 1$ and $2\times 2$ leading principal minors of $Q$ are positive. Let $m_{rs}$ denote the entry of $Q$ at $r$-th row and $s$-th column. Then
@@
Thus the $1\times 1$ leading principal minor of $Q$ is $m_{11}>0$ due to $a_{iik}<0$ and $a_{iij}<0$. And the $2\times 2$ leading principal minor of $Q$ is
@@
Corollary 6.4.
Under the same assumptions as those in Corollary 6.2, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
Proof.
The case (1) follows directly from Theorem 6.3. The proof of the case (2) is analogous to that of Corollary 3.3, and thus we omit the details here. Note that $\Lambda=\sum_{\{ijk\}\in F}\Lambda_{ijk}$, where each $\Lambda_{ijk}$ is derived from either Theorem 5.2 or Theorem 6.3. Q.E.D.
The following theorem establishes the rigidity of the mixed discrete conformal structure III, which generalizes the rigidity part of Theorem 1.8 (vi). As its proof is analogous to that of Theorem 3.4, it is omitted here.
@@
Under the same assumptions as those in Corollary 6.2, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}_{>0}^{N}$.
6.2. Existence of the mixed discrete conformal structure III
The following theorem characterizes the image of $K$ for the mixed discrete conformal structure III, which generalizes the existence part of Theorem 1.8 (vi).
@@ 6.2. Existence of the mixed discrete conformal structure III
Under the same assumptions as Corollary 6.2, the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
To prove Theorem 6.6, we need the following three lemmas.
@@
([5], Lemma 4.6) Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (7). Then $\lim_{f_{j}\rightarrow+\infty}\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$, and the converge is uniform.
Lemma 6.8.
Under the same assumptions as those in Theorem 6.1, if one of the following conditions is satisfied
@@
$\lim f_{i}=-\infty,\ \lim f_{j}=c_{5},\ \lim f_{k}=c_{6}$,
where $c_{1},...,c_{6}\in\mathbb{R}$ are constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
@@
For the case (i), by (48), we derive
(50) $\cosh l_{ij}$ $=e^{f_{i}+f_{j}}(\eta_{ij}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{j}}),$
$\cosh l_{ik}$ $=e^{f_{i}+f_{k}}(\eta_{ik}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{k}}),$
@@
Since $\lim f_{i}=+\infty$, $\lim f_{j}=+\infty$, and $\lim f_{k}=+\infty$, it follows that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$, $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law, we have
@@
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs for the cases (ii) and (iii) follow analogously and are omitted here.
For the case (iv), by (48), we derive
(51) $\cosh l_{ij}=e^{f_{j}-f_{i}}(\eta_{ij}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{j}}+\frac{1}{2}),$
$\cosh l_{ik}=e^{f_{k}-f_{i}}(\eta_{ik}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{k}}+\frac{1}{2}),$
@@
Since $\lim f_{i}=-\infty$, $\lim f_{j}=+\infty$, and $\lim f_{k}=+\infty$, it follows that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$, $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$, and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law, we have
@@
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs for the cases (ii) and (iii) follow analogously and are omitted here. Q.E.D.
@@
Under the same assumptions as those in Theorem 6.1, if one of the following conditions is satisfied
@@
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{4},\ \lim f_{k}=c_{5}$,
where $c_{1},...,c_{5}\in\mathbb{R}$ are constants. Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
@@
For the case (i), it follows from (50) that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law, we have
@@
which implies $\theta^{ik}_{j}\rightarrow 0$. The proofs for the cases (ii), (v) and (vi) follow analogously and are omitted here.
For the case (iii), it follows from (51) that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$$\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$, and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law, we have
@@
which implies $\theta^{ik}_{j}\rightarrow 0$. The proof for the case (iv) follow analogously and is omitted here.
Q.E.D.
Proof of Theorem 6.6: Let $X$ be the image of $K$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows directly from the smooth injectivity of the map. To establish the closedness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that every sequence $\{f^{(m)}\}\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ with $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$ admits a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
Assume, for the sake of contradiction, that there exists a subsequence $\{f^{(m)}\}$ such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$. The boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
@@
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$;
@@
There exists $i\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), such that $f_{i}^{(m)}\rightarrow\pm\infty$;